# Hope vs. Hate: Understanding User Interactions with LGBTQ+ News Content in Mainstream US News Media through the Lens of Hope Speech

### v1

@@ 4 Classification Results
We consider Mistral (Mistral-7B-Instruct-v0.3) [43] and Llama 3 (Llama-3-8B-Instruct) [44], two well-known open LLMs for fine-tuning. For each LLM, we construct four models - three individual political affiliation specific models, and one aggregate model trained on majority label. We determine the aggregate label by taking a majority vote, a standard approach to resolving annotator disagreement (see, e.g., [45, 46]). Any comments that received three separate labels are not used in the aggregate data, accounting for 11.3% of instances, however we used these instances for political affiliation specific models.
@@
Table 3 summarizes our models’ performance on $\mathcal{D}_{\textit{eval}}$. We consider both zero-shot and fine-tuned models (Appendix contains prompt details in § B). For fine-tuning, we use a sequence classification head. We split the training data, after removing the validation set, into an 80/20 test/train split and train for 5 epochs, considering the best model as the one with the highest Macro-F1. Initially, we only consider fine-tuning a model on aggregate labels, but once we determined the highest-performing model, we fine-tuned additional ones specific to each political affiliation (Table 6 describes the noise audit results of these models). We use standard machine learning performance metrics such as accuracy, macro-averaged precision, recall, and F1, as well as the class-specific F1 scores. We find (1) the fine-tuned models considerably outperform zero-shot models; (2) the models’ performance improvement gradually across different stages of the active learning pipeline; and (3) the best-performing model is Llama 3 trained on $\mathcal{D}_{\textit{margin}}$.
Llama [44] Mistral [43]
Accuracy 0.5520 0.8304±0.01 0.8376±0.01 0.8224±0.02 0.6640 0.8024±0.03 0.7640±0.08 0.7840±0.09
Macro-averaged metrics:
Precision 0.5468 0.7286±0.03 0.7378±0.03 0.7127±0.04 0.5654 0.6549±0.12 0.6549±0.07 0.6510±0.12
Recall 0.5547 0.6289±0.01 0.6688±0.02 0.6740±0.01 0.5488 0.5845±0.07 0.6453±0.06 0.6276±0.10
F1 Score 0.4768 0.6507±0.02 0.6838±0.02 0.6850±0.02 0.5375 0.5923±0.08 0.6376±0.07 0.6258±0.11
Class-specific F1 scores:
Positive 0.6845 0.8571±0.01 0.8595±0.01 0.8422±0.02 0.8312 0.8345±0.03 0.7898±0.06 0.7986±0.09
Negative 0.3009 0.5801±0.02 0.6344±0.06 0.6130±0.05 0.4545 0.4269±0.19 0.5620±0.11 0.5487±0.20
Neutral 0.2553 0.2373±0.09 0.3004±0.09 0.3574±0.09 0.1765 0.1971±0.11 0.3037±0.08 0.2381±0.13
Irrelevant 0.6667 0.9281±0.01 0.9409±0.01 0.9276±0.01 0.6879 0.9108±0.02 0.8949±0.08 0.9178±0.04
Table 3: Comparison of Llama and Mistral model results on aggregate labels over five independent training runs. Values are shown as mean±standard deviation where applicable. $\mathcal{M}(.)$ denotes the zero-shot performance. $\mathcal{M}(\mathcal{D})$ denotes model performance when fine-tuned on $\mathcal{D}$. Best results for each metric across all models and settings are shown in bold.
@@ 5 Substantive Findings
We break down our findings into three separate subsections.
5.1 User Engagement with $\mathcal{V}_{\textit{LGBTQ+}}$ Videos
Recall that, $\mathcal{V}_{\textit{LGBTQ+}}$ represents a set of 3,161 news videos relevant to LGBTQ+ issues. In order to compare and contrast user engagement with these videos, we construct a control set (denoted by $\mathcal{V}_{\textit{control}}$). Each video v in $\mathcal{V}_{\textit{LGBTQ+}}$ is matched with a video that is (1) not relevant to LGBTQ+ issues; (2) is uploaded by the same news outlet; (3) has roughly same video duration as v; and (4) has the publishing time closest to v.
Let for a given video $v$, $v_{\textit{likes}}$, $v_{\textit{dislikes}}$, $v_{\textit{views}}$, and $v_{\textit{comments}}$ denote the total number of likes, dislikes, views and comments on $v$, respectively. Following [47, 48, 49, 35], we calculate engagement rate and dislike ratio for a given video as follows.
@@ 5.1 User Engagement with $\mathcal{V}_{\textit{LGBTQ+}}$ Videos
dislikeRate(v) = $\frac{v_{\textit{dislikes}}}{v_{\textit{likes}}+v_{\textit{dislikes}}}$
Table 4 compares the engagement rate and dislike rates of videos in $\mathcal{V}_{\textit{LGBTQ+}}$ and $\mathcal{V}_{\textit{control}}$. YouTube stopped allowing users to access dislikes starting in 2021, so our dislike rate data only includes videos up until that point222sourced from https://github.com/Suji04/YouTube-Dislike-Dataset. We observe that LGBTQ+ videos consistently received higher dislike rates across all three channels compared to non-LGBTQ+ content. This difference is most pronounced for FOX News and CNN, with smaller but still noticeable differences for MSNBC. Overall engagement rates for LGBTQ+ videos are generally lower than for non-LGBTQ+ content, though the magnitude of this difference varied by news outlet. FOX News showed the largest decrease in engagement for LGBTQ+ content, while CNN and MSNBC exhibited differences small enough to likely be insignificant.
The higher dislike ratios and lower engagement rate for LGBTQ+ content across all news outlets, regardless of perceived political leaning, point to a persistent societal resistance to LGBTQ+ topics. Although, this resistance is shown to be more pronounced for the most conservative channel, FOX News.
@@ 5.2 Annotation Study Findings
Prior literature reports association of raters’ political leanings with the way they annotate toxic content [20, 21]. We observe similar results when looking at the agreement between annotators for our task. In Table 5, we observe partisan differences in annotation. The agreement between Democrats and Independents is higher than the agreement between either group and Republicans. This suggests that even given the same comment text and annotation guidelines, Republicans and Democrats or Independents may have different conclusions as to what is positive or negative speech towards the LGBTQ+ community. Even if we collapse the fine-grained labels into HopeSpeech and not-HopeSpeech, the qualitative findings on political disagreement remain unchanged (see, Appendix).
Our survey also include a question asking the raters if they considered themselves a member of the LGBTQ+ (raters can opt out of answering this question), allowing us to analyze if differences in sexual orientation would impact annotation. Out of 375 study participants, 98 identified themselves as being LGBTQ+. The Cohen’s $\kappa$ agreement between LGBTQ+ and non-LGBTQ+ annotators is 0.449. This agreement fits nicely between the low value found when comparing Republicans to others, and the high value found comparing Independents and Democrats. It is quite close to the overall Fleiss’ $\kappa$ of 0.431. This is particularly interesting and indicates that even for a topic highly relevant to gender and sexual orientation, rater disagreement is impacted more by political leanings than sexual orientation. Once again, if we collapse the labels into HopeSpeech and not-HopeSpeech, the qualitative findings on disagreement remain unchanged (see, Appendix).
@@
In addition to our main model fine-tuned on the aggregate labels, we also fine-tune models specifically on the labels of each political affiliation (denoted by $\mathcal{M}_{\textit{rep}}$, $\mathcal{M}_{\textit{dem}}$, and $\mathcal{M}_{\textit{ind}}$). Using each of these models, we classify a set of 50k unseen comments from each channel. Following [50, 21], we conduct a noise audit of these models as shown in Table 6. While we observe the in-the-wild agreement across models is higher than the agreement across human raters with different political beliefs, the impact of the political affiliations still shone through. Once again, Democrats and Independents share higher agreement with each other, while they both have lower agreement with Republicans to a similar degree. This demonstrates that the human biases due to political affiliation in labelled data can perpetuate to fine-tuned models. Tables 78 present illustrative examples showing how trans-exclusionary stance of human Republican raters may get perpetuated to fine-tuned LLM $\mathcal{M}_{\textit{rep}}$ affecting marginalized voices.
$\mathcal{M}_{\textit{dem}}$ 0.572 0.670
$\mathcal{M}_{\textit{rep}}$ 0.572 - 0.573
$\mathcal{M}_{\textit{ind}}$ 0.670 0.573 -
damn! first time I’ve seen Tucker actually lose. I love you Tucker, but the military transgender ban is BS. it’s not a huge expense. Like Sean Patrick said in the end before Tucker rage quits "They’re earning those benefits by putting their life on the line for America and have done more for this country than you (Tucker) and I ever will". Hormones arent expensive. 10ml testosterone costs \$40 and it lasts them 6 months. Negative Positive
Table 7: Illustrative examples were $\mathcal{M}_{\textit{rep}}$ does not agree with $\mathcal{M}_{\textit{dem}}$ and $\mathcal{M}_{\textit{ind}}$.
@@
We also calculate the performance metrics of zero-shot classification against the annotations of each political affiliation (see, Table 9). The results show similar performance with Democrat and Independent labels, but each metrics was lower for the Republican labels. This was present across the three different models we tested; GPT-4o-mini, Llama, and Mistral. This suggests that the biases present in these models may align closer with the biases found in our human Independent and Democrat annotators, than those of a Republican.
@@
5.3 In-the-wild Hope Speech Analyses
@@
In order to look at the hope and hate speech make-up of LGBTQ+ discussion in-the-wild, we consider our best-performing model and classify 50k randomly sampled, unseen comments from each news outlet. The results, as shown in Figure 2, reveal distinct patterns. Notably, FOX News stands out with the highest proportion of negative comments (24.15%) and the lowest proportion of positive comments (2.3%) among the three outlets. This aligns with the outlet’s reputation for conservative viewpoints. In contrast, CNN shows a more balanced distribution, with the highest percentage of positive comments (7.66%) and a negative comment rate (17.2%) closer to the overall average. MSNBC, interestingly, has a much lower proportion of positive comments than CNN (4.55%) and the lowest proportion of negative comments (9.36%).
To get a better understanding of the sentiment of these comments sections we introduce the Positivity Ratio.
Positivity Ratio=#Positive Comments#Positive Comments+#Negative $\text{Positivity Ratio}=\frac{\#\text{Positive Comments}}{\#\text{Positive Comments}+\#\text{Negative Comments}}$
IrrelevantNegativePositiveNeutral0020202020404040406060606080808080100100100100PercentageOverallCNNFOXMSNBC
Figure 2: Label breakdown by news outlets with 95% confidence intervals. 50k comments from each news outlet found in-the-wild were classified by our best performing model.
Label Overall CNN FOX MSNBC
Negative 16.90% ± 0.025% 17.20% ± 0.05% 24.15% ± 0.045% 9.36% ± 0.04%
Positive 4.84% ± 0.015% 7.66% ± 0.035% 2.30% ± 0.02% 4.55% ± 0.02%
Neutral 1.73% ± 0.01% 2.19% ± 0.02% 1.77% ± 0.02% 1.22% ± 0.01%
Positive Ratio 22.25% ± 0.065% 30.81% ± 0.12% 8.70% ± 0.07% 32.70% ± 0.15%
Table 10: Label breakdown and positivity rates by channel with 95% confidence intervals. 50k comments from each channel found in-the-wild were classified by our best performing model.
@@
This ratio focuses solely on comments expressing clear sentiment, excluding irrelevant and neutral comments to provide a more direct measure of sentiment polarity. By using this metric, we find that MSNBC has the highest positivity ratio at 32.7%, followed closely by CNN at 30.81%, both above the overall average of 22.25%. In stark contrast, FOX News shows a significantly lower positivity ratio of 8.7%. These findings further emphasize the divergence in audience sentiment across channels, with MSNBC and CNN fostering more balanced or slightly positive discussions around LGBTQ+ content, while FOX News comments lean heavily towards negative sentiment.
@@
Perhaps the most important results come from looking at the trends as a whole. For every channel, there were more negative comments than positive ones. This consistent imbalance, favoring negative sentiment across all channels, suggests a broader societal tendency towards critical or oppositional engagement with LGBTQ+ topics in online spaces. It implies that, regardless of the platform, viewers are more inclined to express disapproval or criticism than support or affirmation when commenting on LGBTQ+-related content. Another noteworthy trend is the uniformly low percentage of neutral comments across all channels, ranging from 1.22% to 2.22%. This scarcity of neutral perspectives, coupled with the positive-negative imbalance, points to a highly polarized discourse surrounding LGBTQ+ issues. It suggests that those who engage in these comment sections tend to hold and express strong opinions, whether positive or negative, rather than maintaining a neutral stance.
5.4 Error Analysis
While our model performed well in identifying the rare class of LGBTQ+ hope speech in the wild, there were cases where the beliefs expressed in comments were too nuanced to be classified correctly. Table 11 showcases two such examples. Both comments are generally supportive on a surface level. One is promoting human rights and social progress, the other is giving respect to female athletes. However, in context it is clear the support is not directed to the LGBTQ+. The first comment is equating the social progress of the LGBTQ+ community with bestiality, a common homophobic cliché. The second comment is specifically singling out "biological female athletes" to support, taking a clear stance on the issue of transgender people in sports.
Examples like these illustrate the challenges a task like this entails. Our model fails to fully understand the sarcasm or superficial support that less straight-forward comments may contain. Enhancing the model’s ability to understand context and implicit meanings, possibly through the use of more advanced language models or improved tuning, is a worthy future research challenge.
YouTube Comment Model Prediction
Table 11: Two example challenging comments our classifier mistakenly identified as Positive.

### v2

@@ 4 Classification Results
We consider Mistral (Mistral-7B-Instruct-v0.3) Jiang et al. (2023) and Llama 3 (Llama-3-8B-Instruct) Dubey et al. (2024), two well-known open LLMs for fine-tuning. We determine the aggregate label by taking a majority vote, a standard approach to resolving annotator disagreement (see, e.g., Davidson et al., 2017; Wiegand et al., 2019). Any comments that received three separate labels are not used in the aggregate data, accounting for 11.3% of instances. In addition to fine-tuning models on aggregate labels, we also fine-tune models specific to each political affiliation. For those models, we use these 11.3% additional instances [^3] Table 8 describes the noise audit results of these models..
@@
Table 3 summarizes our models’ performance on $\mathcal{D}_{\textit{eval}}$. Collectively, we refer to our annotated corpus as $\mathcal{D}_{\textit{hope}}$, which evolved through three active learning phases as seen in Table 2: initial seed collection ($\mathcal{D}_{\textit{seed}}$) comprising 1,950 instances, certainty sampling ($\mathcal{D}_{\textit{certainty}}$) which expanded the dataset to 2,850 instances, and margin sampling ($\mathcal{D}_{\textit{margin}}$) which yielded our final dataset of 3,750 instances. $\mathcal{D}_{\textit{eval}}$ consists of 250 instances randomly selected from $\mathcal{D}_{\textit{seed}}$ and held out from all training processes, serving as a consistent evaluation set for all models. We consider both zero-shot and fine-tuned models (Appendix contains prompt details in § D). For fine-tuning, we use a sequence classification head. We split the training data, after removing the validation set, into an 80/20 test/train split and train for 5 epochs, considering the best model as the one with the highest Macro-F1. We find (1) the fine-tuned models considerably outperform zero-shot models; (2) the models’ performance improve gradually across different stages of the active learning pipeline; and (3) the best-performing model is Llama 3 trained on $\mathcal{D}_{\textit{margin}}$.
For the sake of completeness, we compare performance against several off‐the‐shelf classifiers on adjacent tasks such as hate speech detection, sentiment classification, and stance detection (results in Table A1). As sentiment classifier baselines, we consider a DistilBERT model Sanh et al. (2019) fine-tuned on SST-2 Socher et al. (2013), which provides only positive/negative predictions. For fair comparison, we map positive sentiment to hope speech and negative sentiment to negative content, while removing any comments with neutral or irrelevant content. It achieves a Macro F1 of 0.601 and Positive F1 of 0.520, significantly lower than our best Llama-based model (Macro F1: 0.864, Positive F1: 0.685), even with the removal of neutral/irrelevant label noise. This gap stems from sentiment classifiers failing to capture intent in LGBTQ+ discourse. For instance, a sentiment classifier will predict I will be extremely happy if Congress passes a resolution and takes away voting rights from the LGBTQ+ people as positive.
We further compare our approach with toxicity and hate speech detectors. We use Detoxify Hanu and Unitary team (2020) for toxic comment detection and cardiffnlp’s hate-roBERTa Barbieri et al. (2020) for hate speech detection. For fair comparison, we map toxic and hateful content to negative speech and hope, neutral, or irrelevant content to non–toxic and non-hateful speech. Therefore, we report their Negative F1 scores (i.e., their ability to correctly identify harmful content) rather than Positive F1. These models achieved low Negative F1 scores (0.129 and 0.258 respectively), as they focus on overt toxicity, while in LGBTQ+ discourse disapproval can often manifest without explicit hate speech.
Finally, our task bears many similarities to that of determining stance on the LGBTQ+ community. However, off-the-shelf stance detectors are typically developed for specific topics (e.g., abortion or climate change), and there are few systems available that allow for input of a custom topic beyond using LLMs to specify in the prompt. In practice, these LLM-based approaches for stance detection end up resembling the zero-shot evaluations we have already reported.
Error analysis. Consistent with the existing literature Joshi et al. (2017); Farabi et al. (2024), we observe that our models often struggled in detecting sarcasm. Appendix (§ N) discusses challenging examples.
@@ 5 Substantive Findings
5.1 User Engagement Findings
$\mathcal{V}_{\textit{LGBTQ+}}$ represents a set of 3,161 news videos relevant to LGBTQ+ issues. We first construct a control set, $\mathcal{V}_{\textit{control}}$. Each video v in $\mathcal{V}_{\textit{LGBTQ+}}$ is matched with a video that is (1) not relevant to LGBTQ+ issues; (2) is uploaded by the same news outlet; (3) has roughly same video duration as v; and (4) has the publishing time closest to v.
Let for a given video $v$, $v_{\textit{likes}}$, $_{\textit{dislikes}}, vv$$_{\textit{views}}$, and $v_{\textit{comments}}$ denote the total number of likes, dislikes, views and comments on $v$, respectively. Following López-Navarrete et al., 2021; Le et al., 2022; Braun et al., 2017; KhudaBukhsh et al., 2021, we calculate engagement rate and dislike ratio for a given video as follows.
@@ 5.1 User Engagement Findings
dislikeRate($v$) = $\frac{v_{\textit{dislikes}}}{v_{\textit{dislikes}}+v_{\textit{likes}}}$
@@
Table 4 compares the engagement rate and dislike rates of videos in $\mathcal{V}_{\textit{LGBTQ+}}$ and $\mathcal{V}_{\textit{control}}$. YouTube stopped allowing users to access dislikes starting in 2021, so our dislike rate data only includes videos up until that point444sourced from Dutta et al., 2024b. We observe that LGBTQ+ videos consistently received higher dislike rates across all three channels compared to non-LGBTQ+ content. This difference is most pronounced for FOX News and CNN, with smaller but still noticeable differences for MSNBC. Overall engagement rates for LGBTQ+ videos are generally lower than for non-LGBTQ+ content, though the magnitude of this difference varied by news outlet. FOX News showed the largest decrease in engagement for LGBTQ+ content, while CNN and MSNBC exhibited differences small enough to likely be insignificant. The higher dislike ratios and lower engagement rate for LGBTQ+ content across all news outlets, regardless of perceived political leaning, point to a persistent societal resistance to LGBTQ+ topics. Although, this resistance is shown to be more pronounced for the most conservative channel, FOX News. Appendix K and L contain additional contrastive analyses with other hot-button issues and a deeper look into heavily disliked videos.
@@ 5.2 Annotation Study Findings
Table 5 suggests partisan differences in the annotation of hope speech for the LGBTQ+ community. Even given the same comment text and annotation guidelines, Republicans and Democrats or Independents may have different conclusions as to what is hope speech or negative towards the LGBTQ+ community. The agreement between Democrats and Independents is higher than the agreement between either group and Republicans.
Our survey also includes a question [^5] Raters could opt out of answering this question. asking the raters if they considered themselves a member of the LGBTQ+ community. Out of 375 raters, 98 (31 Democrats; 29 Republicans; and 38 Independents) identified themselves as being LGBTQ+. The Cohen’s $\kappa$ agreement between LGBTQ+ and non-LGBTQ+ annotators is 0.449. This agreement fits between the low values found when comparing Republicans to others, and the high value found comparing Independents and Democrats and is close to the overall Fleiss’ $\kappa$ of 0.431. We find this result particularly interesting as it indicates that even for a topic highly relevant to gender and sexual orientation, rater disagreement is perhaps impacted more by political leanings than sexual orientation. Even if we collapse the labels into HopeSpeech and not-HopeSpeech, the qualitative findings on political disagreement remain unchanged (see, Appendix).
To better understand intra-group annotation agreement, we collect additional annotations for a subset of 30 batches where each instance was labeled by two annotators of the same political affiliation. Independent annotators show the highest internal consistency ($\kappa$ = 0.500), suggesting more unified perspectives within this group than the Democrats ($\kappa$ = 0.446) and the Republicans ($\kappa$ = 0.368). The Republicans show the lowest intra-group agreement (even lower than their inter-group agreement with Democrats or Independents) suggesting more varied perspectives within the group.
To analyze the intersection of LGBTQ+ identity and political beliefs, we next conduct two-proportion z-tests comparing label distributions between LGBTQ+ and non-LGBTQ+ annotators within each political affiliation. The analysis has two key major takeaways (details in Table A17). First, LGBTQ+ annotators across all political affiliations were consistently less likely to label content as irrelevant compared to their non-LGBTQ+ counterparts (Democrats: -4.1%, p < 0.05; Republicans: -6.0%, p < 0.001; Independents: -3.5%, p < 0.05), suggesting they more readily recognize content relevant to LGBTQ+ issues. Table A13 shows a comment that a LGBTQ+ annotator correctly identified as LGBTQ+-related, while the two other annotators seemed to have not enough context to make the connection. Second, the strongest differences appear among Republican annotators, where LGBTQ+ Republicans were significantly more likely to label content as negative and less likely to label it as positive compared to non-LGBTQ+ Republicans. These findings suggest that lived experience as an LGBTQ+ person may affect content interpretation, with the strongest differences emerging for ideologically charged content among conservative annotators.
@@
In addition to our main model fine-tuned on the aggregate labels, we also fine-tune models specifically on the labels of each political affiliation (denoted by $\mathcal{M}_{\textit{rep}}$, $\mathcal{M}_{\textit{dem}}$, and $\mathcal{M}_{\textit{ind}}$). $\mathcal{M}_{\textit{rep}}$ was trained only on the 3,750 comments labeled by Republicans, $\mathcal{M}_{\textit{dem}}$ only on the 3,750 labeled by Democrats, and $\mathcal{M}_{\textit{ind}}$ on the 3,750 labeled by Independents. Using each of these models, we classify a set of 50k unseen comments from each channel. Following Kahneman et al. (2021); Weerasooriya et al. (2023), we conduct a noise audit of these models as shown in Table 8. A noise audit involves examining agreement between different model variants on identical inputs, revealing systematic differences in model predictions. To verify whether political biases are introduced during fine-tuning, we first examine agreement between pre-trained and fine-tuned models. We observe notably low agreement scores: Cohen’s $\kappa$ of 0.29 for both $\mathcal{M}_{\textit{ind}}$ and $\mathcal{M}_{\textit{dem}}$ with pre-trained Llama, and 0.25 for $\mathcal{M}_{\textit{rep}}$ with pre-trained Llama. These low agreement scores suggest substantial shifts in model behavior during fine-tuning. While we observe the in-the-wild agreement across models is higher than the agreement across human raters with different political beliefs, the impact of the political affiliations still shone through. Once again, $\mathcal{M}_{\textit{dem}}$ and $\mathcal{M}_{\textit{ind}}$ share higher agreement with each other, while they both have lower agreement with $\mathcal{M}_{\textit{rep}}$ to a similar degree. This demonstrates that the human biases due to political affiliation in labeled data can perpetuate to fine-tuned models. Tables 6 and 7 present illustrative examples showing how trans-exclusionary stance of human Republican raters may get perpetuated to $\mathcal{M}_{\textit{rep}}$ affecting marginalized voices.
@@
damn! first time I’ve seen Tucker actually lose. I love you Tucker, but the military transgender ban is BS. it’s not a huge expense. Like Sean Patrick said in the end before Tucker rage quits "They’re earning those benefits by putting their life on the line for America and have done more for this country than you (Tucker) and I ever will"… Negative Positive
Table 7: Illustrative examples where $\mathcal{M}_{\textit{rep}}$ does not agree with $\mathcal{M}_{\textit{dem}}$ and $\mathcal{M}_{\textit{ind}}$.
$\mathcal{M}_{\textit{dem}}$ - 0.572 0.670
$\mathcal{M}_{\textit{rep}}$ 0.572 - 0.573
$\mathcal{M}_{\textit{ind}}$ 0.670 0.573 -
We also calculate the performance metrics of zero-shot classification against the annotations of each political affiliation (see, Table 9). The results show similar performance with Democrat and Independent labels, but each metrics was lower for the Republican labels. Similar trend was present across the three different models we tested; GPT-4o-mini, Llama, and Mistral. This suggests that the biases present in these models may align closer with the biases found in our human Independent and Democrat raters, than those of a Republican.
In addition, we conduct a parallel analysis training separate models on annotations from LGBTQ+ and non-LGBTQ+ annotators. For fair comparison, in this experiment, we use a reduced dataset where comments had labels from both LGBTQ+ and non-LGBTQ+ annotators. Consistent with our findings from human annotators, we observe that the non-LGBTQ+ model more frequently classified content as "irrelevant" compared to the LGBTQ+ model. Table A13 shows an example where the LGBTQ+ model identified as relevant content that non-LGBTQ+ annotators marked as irrelevant.
@@
Key takeaways: (1) Annotation of hope speech is associated with political leanings that drive more disagreement than gender and sexual identity on a task relevant to gender and sexual identity; (2) LGBTQ+ annotators across all political affiliations were consistently less likely to label content as irrelevant and showed stronger recognition of LGBTQ+-related content, with the most pronounced differences appearing among Republican annotators; (3) political biases in human annotations carried over to fine-tuned language models, with models trained on Democrats and Independents showing higher agreement compared to the Republican trained model.
5.3 In-the-wild hope speech findings
@@
In order to look at the hope and hate speech make-up of LGBTQ+ discussion in-the-wild, we consider our best-performing model and classify 50k randomly sampled, unseen comments from each news outlet. The results, as shown in Figure 2 with numerical results in Table A19, reveal distinct patterns. In particular, FOX News has the highest proportion of negative comments (24.15%) and the lowest proportion of hope speech (2.3%) among the three outlets. This aligns with the outlet’s reputation for conservative viewpoints. In contrast, CNN shows a more balanced distribution, with the highest percentage of hope speech (7.66%) and a negative comment rate (17.2%) closer to the overall average. MSNBC, however, has a much lower proportion of hope speech than CNN (4.55%) and the lowest proportion of negative comments (9.36%).
@@
Figure 2 reports the ratio of hope speech and hope speech and negatives found in the wild. Dubbed positivity ratio, this ratio focuses on comments expressing polarized stance towards the LGBTQ+ community, excluding irrelevant and neutral comments. We find that MSNBC has the highest positivity ratio at 32.7%, followed by CNN at 30.81%, both above the overall average of 22.25%. In stark contrast, FOX News shows a considerably lower positivity ratio of 8.7%. These findings further emphasize the divergence in audience sentiment across channels, with MSNBC and CNN fostering more balanced or slightly supportive discussions around LGBTQ+ content, while FOX News comments lean heavily towards negative sentiment. To better understand the nature of these discussions beyond labels, Appendix M contains a lexical analysis of the language used in different comment categories.
IrrelevantNegativePositiveNeutral02020404060608080100100PercentageOverallCNNFOXMSNBC
Figure 2: Label breakdown by news outlets with 95% confidence intervals. 50k comments from each news outlet found in-the-wild were classified by our best performing model. Numerical values are listed in Table A19.
@@
Perhaps the most important results come from looking at the trends as a whole. For every channel, there were more negative comments than hope speech. This consistent imbalance, suggests a broader societal tendency towards critical or oppositional engagement with LGBTQ+ topics in online spaces. It implies that, regardless of the audience, viewers are more inclined to express disapproval or criticism than support or affirmation when commenting on LGBTQ+-related content.  [^6] Our Limitations and Appendix contain an exploratory study into Reddit showing that there are certain social web pockets which may be more supportive than negative, warranting additional investigation into more varied social media. Another noteworthy trend is the uniformly low percentage of neutral comments across all channels, ranging from 1.22% to 2.22%. This scarcity of neutral perspectives, coupled with the positive-negative imbalance, points to a highly polarized discourse surrounding LGBTQ+ issues. It suggests that those who engage in these comment sections tend to hold and express strong opinions, whether positive or negative, rather than maintaining a neutral stance.
To confirm the accuracy of our model when generalizing to in-the-wild data, we conducted a manual inspection of 100 comments labelled as Positive. From those 100, we found 16 false positives, mainly consisting of comments that we would label as Negative according to our guidelines, such as:
LGBTQHDJFNEIEKBEJDJDNDIDBDJIDJDIGWJOQPQPLAIDHWNRBHXKSOFORBRIDKWNDIKFKDNRUWODUE Get it together bigot!
This example shows a comment that is Negative towards the LGBTQ+ community in a more subtle or sarcastic way. An error analysis in Appendix N further discusses some limitations that we identified through challenging comments it labeled incorrectly.